Calculate how many moles of NH3 form when each quantity of reactant completely reacts according to the equation: 3N2H4(l)→4NH3(g)+N2(g)
5.2 mol N2H4
2.55 mol N2H4
36.3 g N2H4
14.8 kg N2H4
1
Expert's answer
2017-03-09T09:38:06-0500
n(NH_3 )=4/3 n(N_2 H_4 )
n_1 (NH_3 )=4/3×5.2 mol=6.9 mol
n_2 (NH_3 )=4/3×2.55 mol=3.40 mol
n_3 (NH_3 )=4/3×36.3 mol=48.4 mol
n_4 (NH_3 )=4/3×14.8 mol=19.7 mol